# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ Appendix A Moore determinants
In this section, we collect several results on Moore determinants. With the exception of Lemma A.6, for which we provide a proof, these results are already known (from work of Elkies in [13] and Fresnel-Matignon in [10]), and are therefore recalled without proof.
@@
Moreover, given an invertible matrix $M\in GL_{n}({\mathbb{F}}_{p})$, we have ([13, p. 80]) that
@@
For every $n$-tuple $\underline{a}\in k^{n}$, and for every $1\leq i\leq n$, we define the $n-1$-tuple $\underline{\hat{a_{i}}}:=(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n})$. Then, we have the formula
@@
This is the special case $m=0$ of [10, Theorem 4.1]. ∎
@@
Lemma A.5.
(cf. Proposition 2.2. of [10]) Let $V\subset k$ be a ${\mathbb{F}}_{p}$-vector space of dimension $n$. The structural polynomial $P_{V}$ is additive and for every choice of basis $\underline{v}=(v_{1},\dots,v_{n})$ of $V$ it satisfies the identity
@@
Let $n\geq 1,\ m\geq 1$ and $Y_{1},...,Y_{n}$, $X_{1},...,X_{m}$ be variables over ${\mathbb{F}}_{p}$. Then
$\Delta_{n}(\Delta_{m+1}(Y_{1},X_{1},X_{2},...,X_{m}),...,\Delta_{m+1}(Y_{n},X_{1},X_{2},...,X_{m}))=$
(A.7) $=\Delta_{m}(X_{1},...,X_{m})^{p+p^{2}+...+p^{n-1}}\Delta_{n+m}(Y_{1},...,Y_{n},X_{1},...X_{m}).$
@@
We proceed by induction on $n$. If $n=1$, we interpret the expression $p+p^{2}+...+p^{n-1}$ as 0. If $n=2$, then the Lemma is a special case of Theorem A.3. Let us then assume $n>2$ and that the Lemma is satisfied when replacing $n$ with $n-1$.
We denote by $\mathcal{X}$, the ${\mathbb{F}}_{p}$-vector space $\langle X_{1},X_{2},\dots,X_{m}\rangle_{{\mathbb{F}_{p}}}$, and by $P_{\mathcal{X}}$ the structural polynomial of $\mathcal{X}$. It then follows from Lemma A.5 that
$\Delta_{n}(\Delta_{m+1}(Y_{1},X_{1},X_{2},...,X_{m}),...,\Delta_{m+1}(Y_{n},X_{1},X_{2},...,X_{m}))=$
$\Delta_{n}((-1)^{m}\Delta_{m}(\underline{X})P_{\mathcal{X}}(Y_{1}),\cdots,(-1)^{m}\Delta_{m}(\underline{X})P_{\mathcal{X}}(Y_{n}))=$
$(-1)^{mn}\Delta_{m}(\underline{X})^{1+p+\cdots+p^{n-1}}\Delta_{n}(P_{\mathcal{X}}(Y_{1}),\cdots,P_{\mathcal{X}}(Y_{n})).$
@@
(A.8) $\Delta_{n+m}(Y_{1},...,Y_{n},X_{1},...X_{m})=(-1)^{mn}\Delta_{m}(\underline{X})\Delta_{n}(P_{\mathcal{X}}(Y_{1}),\cdots,P_{\mathcal{X}}(Y_{n})).$
@@
$\Delta_{n+m}(Y_{1},...,Y_{n},X_{1},...X_{m})=(-1)^{m}\Delta_{n+m}(Y_{1},...,Y_{n-1},X_{1},...X_{m},Y_{n})=$
$(-1)^{m}\Delta_{n-1+m}(Y_{1},...,Y_{n-1},X_{1},..,X_{m})\prod_{v\in\langle Y_{1},..,Y_{n-1},X_{1},..,X_{m}\rangle_{{\mathbb{F}_{p}}}}(Y_{n}+v),$
and then we can apply the identity (A.8) replacing $n$ with $n-1$ (satisfied by the inductive hypothesis) to show that (A.7) is equivalent to
$\Delta_{n}(P_{\mathcal{X}}(Y_{1}),\cdots,P_{\mathcal{X}}(Y_{n}))=\Delta_{n-1}(P_{\mathcal{X}}(Y_{1}),\cdots,P_{\mathcal{X}}(Y_{n-1}))\prod_{v\in\langle Y_{1},..,Y_{n-1},X_{1},..,X_{m}\rangle}(Y_{n}+v)$
@@
$\prod_{w\in\langle P_{\mathcal{X}}(Y_{1}),\cdots,P_{\mathcal{X}}(Y_{n-1})\rangle_{{\mathbb{F}_{p}}}}(P_{\mathcal{X}}(Y_{n})+w)=\prod_{v\in\langle Y_{1},..,Y_{n-1},X_{1},..,X_{m}\rangle_{{\mathbb{F}_{p}}}}(Y_{n}+v).$
@@
When $m=1$ the identity of Lemma A.6 is already known. It appears for example in work of Elkies [13, p.81] and it is used in the proof of Proposition 5.3.
@@
Proposition A.10.
(Proposition 2.3. of [10]) Let $W\subset k$ be a ${\mathbb{F}}_{p}$ vector space of dimension $n$. Let $\underline{a}$ be an ${\mathbb{F}}_{p}$-basis of $W$ and $\underline{a}^{\star}=(a_{1}^{\star},\dots,a_{n}^{\star})$ be its dual basis. For every non-zero $n$-tuple $\underline{\epsilon}=(\epsilon_{1},\dots,\epsilon_{n})\in{\mathbb{F}}_{p}^{n}-\underline{0}$, we denote by $\varphi_{\underline{\epsilon}}$ the linear homomorphism $W\to{\mathbb{F}}_{p}$ given by $\varphi_{\underline{\epsilon}}:=\sum_{i=1}^{n}\epsilon_{i}a_{i}^{\star}$. Recall that we have defined $P_{\ker\varphi_{\underline{\epsilon}}}(X):=\prod_{w\in\ker\varphi_{\underline{\epsilon}}}(X-w)$.
@@
(A.14) $\delta_{\underline{\epsilon}}(\underline{a})=\sum_{i=1}^{n}(-1)^{i-1}\epsilon_{i}\Delta_{n-1}(\underline{\hat{a_{i}}}).$
Proposition A.15 (Proposition 5.1. of [10]).
@@
Finally, we prove two results that are not in [10] and that are used in Section 4.
@@
Let $X_{1},...,X_{n}$ be variables over ${\mathbb{F}_{p}}$, and for every $0\leq t\leq n$ let $V_{n-t}=\langle X_{1},...,X_{n-t}\rangle_{{\mathbb{F}}_{p}}$. Then, we have that
$\Delta_{n}(X_{1},...,X_{n})=\Delta_{n-t}(X_{1},...,X_{n-t})\Delta_{t}(P_{V_{n-t}}(X_{n-t+1}),...,P_{V_{n-t}}(X_{n})),$
@@
$\frac{\Delta_{n-1}(X_{1},..,\widehat{X_{i}},...,X_{n})}{\Delta_{n}(X_{1},...,X_{n})}=\frac{\Delta_{t-1}(P_{V_{n-t}}(X_{n-t+1}),...,\widehat{P_{V_{n-t}}(X_{i})},...,P_{V_{n-t}}(X_{n}))}{\Delta_{t}(P_{V_{n-t}}(X_{n-t+1}),...,P_{V_{n-t}}(X_{n}))}.$
@@
Let $n\geq 1$, let $M\in\operatorname{GL}_{n}({\mathbb{F}}_{p})$ and let $X_{1},...,X_{n}$ be variables over ${\mathbb{F}}_{p}$. Denote by $Y_{i}=\Delta_{n-1}(\hat{\underline{X}_{i}})$ and by $Y_{i}^{M}=\Delta_{n-1}\left(\widehat{(\underline{X}M)_{i}}\right)$. Then, we have

### v2

@@ Appendix A Moore determinants and invariant theory
In this section, we collect several results on Moore determinants. With the exception of Lemma A.7 and Lemma A.19, for which we provide a proof, these results are already known (from work of Dickson [Dickson11], Elkies in [Elkies99] and Fresnel-Matignon in [FresnelMatignon23]), and are therefore recalled with minimal or no proof.
@@
Moreover, given an invertible matrix $M\in GL_{n}({\mathbb{F}}_{p})$, we have ([Elkies99, p. 80]) that
@@
For every $n$-tuple $\underline{a}\in k^{n}$, and for every $1\leq i\leq n$, we define the n1n-1-tuple $\underline{\hat{a_{i}}}:=(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{n})$. Then, we have the formula
@@
This is the special case $m=0$ of [FresnelMatignon23, Theorem 4.1]. ∎
@@
Lemma A.5 (Proposition 2.2. of [FresnelMatignon23]).
Let $V\subset k$ be a ${\mathbb{F}}_{p}$-vector space of dimension $n$. The structural polynomial $P_{V}$ is additive and for every choice of basis $\underline{v}=(v_{1},\dots,v_{n})$ of $V$ it satisfies the identity
@@
The structural polynomial is tightly related to the invariant theory of finite groups, via its tight connections to Dickson invariants: if ${\mathbb{F}}_{p}[X_{1},\dots,X_{n}]$ is a polynomial ring on $n$ variables over ${\mathbb{F}}_{p}$, and $V_{n}$ is the ${\mathbb{F}}_{p}$-vector space $\langle X_{1},\dots,X_{n}\rangle_{{\mathbb{F}_{p}}}$, one defines the Dickson invariants $c_{n,i}\in{\mathbb{F}}_{p}[X_{1},\dots,X_{n}]$ for $0\leq i\leq n$ as the coefficients of the structural polynomial of $V_{n}$. More precisely, we have
$P_{V_{n}}(T)=\frac{\Delta_{n+1}(\underline{X},T)}{\Delta_{n}(\underline{X})}=\prod_{v\in V_{n}}(T-v)=\sum_{i=0}^{i=n}(-1)^{n-i}c_{n,i}T^{p^{i}}\in{\mathbb{F}}_{p}[X_{1},\dots,X_{n}][T].$
Note that the $c_{n,i}$’s are invariant under the natural action of $\operatorname{GL}_{n}({\mathbb{F}}_{p})$ on ${\mathbb{F}}_{p}[X_{1},\dots,X_{n}]$ and Dickson proved that they generate the algebra of polynomials that are invariant by this action. In the following proposition, we collect the properties of Dickson invariants that will be used in Section 4.2.
Proposition A.6.
Let ${\mathbb{F}}_{p}[X_{1},\dots,X_{n}]$ be the polynomial ring on $n$ variables over ${\mathbb{F}}_{p}$, and for every $1\leq j\leq n$ let $P_{V_{j}}$ be the structural polynomial of the ${\mathbb{F}}_{p}$-vector space $V_{j}=\langle X_{1},\dots,X_{j}\rangle_{{\mathbb{F}_{p}}}$. For $0\leq i<j\leq n$ denote by $c_{j,i}$ the $i$-th Dickson invariant in the subring ${\mathbb{F}}_{p}[X_{1},\dots,X_{j}]$. Then, we have the following identities
  1. (i)
    $P_{V_{n-1}}(X_{n})^{p-1}=\left(\frac{\Delta_{n}(X_{1},\dots,X_{n})}{\Delta_{n-1}(X_{1},\dots,X_{n-1})}\right)^{p-1}=c_{n,n-1}-c_{n-1,n-2}^{p}.$
  2. (ii)
    $c_{n,i}=\frac{\det(\underline{X},\underline{X}^{p},\cdots,\widehat{\underline{X}^{p^{i}}},\cdots,{\underline{X}}^{p^{n}})}{\Delta_{n}(X_{1},\dots,X_{n})}.$
  3. (iii)
    $c_{n,n-1}=\sum_{\lceil\underline{x}\rceil\in\mathbb{P}({\mathbb{F}}_{p}^{n})}\ \prod_{\lceil\underline{\epsilon}\rceil\in\mathbb{P}({\mathbb{F}}_{p}^{n})|\sum_{i}\epsilon_{i}x_{i}\neq 0}(\sum_{i}\epsilon_{i}X_{i})^{p-1}.$
Proof.
  1. (i)
    See [Wilkerson83, Proposition 1.3 b)]
  2. (ii)
    See [Wilkerson83, Proposition 1.3 a)]
  3. (iii)
    This is a special case of a theorem by Stong and Tamagawa as stated in [NeuselSmith10, Theorem 6.1.7.]
@@
Let $n\geq 1,\ m\geq 1$ and $Y_{1},\dots,Y_{n}$, $X_{1},\dots,X_{m}$ be variables over ${\mathbb{F}}_{p}$. Then
$\Delta_{n}(\Delta_{m+1}(Y_{1},X_{1},X_{2},\dots,X_{m}),\dots,\Delta_{m+1}(Y_{n},X_{1},X_{2},\dots,X_{m}))=$
(A.8) $=\Delta_{m}(X_{1},\dots,X_{m})^{p+p^{2}+\dots+p^{n-1}}\Delta_{n+m}(Y_{1},\dots,Y_{n},X_{1},\dots,X_{m}).$
@@
We proceed by induction on $n$. If $n=1$, we interpret the expression $p+p^{2}+\dots+p^{n-1}$ as 0. If $n=2$, then the Lemma coincides with [FresnelMatignon23, Corollary 4.2] (also specialized to the case $n=2$ in the notation of the Corollary). Let us then assume $n>2$ and that the Lemma is satisfied when replacing $n$ with n1n-1.
We denote by $V_{m}$, the ${\mathbb{F}}_{p}$-vector space $\langle X_{1},X_{2},\dots,X_{m}\rangle_{{\mathbb{F}_{p}}}$, and by $P_{V_{m}}$ the structural polynomial of $V_{m}$. It then follows from Lemma A.5 that
$\Delta_{n}(\Delta_{m+1}(Y_{1},X_{1},X_{2},\dots,X_{m}),\dots,\Delta_{m+1}(Y_{n},X_{1},X_{2},\dots,X_{m}))=$
$\Delta_{n}((-1)^{m}\Delta_{m}(\underline{X})P_{V_{m}}(Y_{1}),\cdots,(-1)^{m}\Delta_{m}(\underline{X})P_{V_{m}}(Y_{n}))=$
$(-1)^{mn}\Delta_{m}(\underline{X})^{1+p+\cdots+p^{n-1}}\Delta_{n}(P_{V_{m}}(Y_{1}),\cdots,P_{V_{m}}(Y_{n})).$
@@
(A.9) $\Delta_{n+m}(Y_{1},\dots,Y_{n},X_{1},\dots,X_{m})=(-1)^{mn}\Delta_{m}(\underline{X})\Delta_{n}(P_{V_{m}}(Y_{1}),\cdots,P_{V_{m}}(Y_{n})).$
@@
$\Delta_{n+m}(Y_{1},\dots,Y_{n},X_{1},\dots,X_{m})=(-1)^{m}\Delta_{n+m}(Y_{1},\dots,Y_{n-1},X_{1},\dots,X_{m},Y_{n})=$
$(-1)^{m}\Delta_{n-1+m}(Y_{1},\dots,Y_{n-1},X_{1},\dots,X_{m})\prod_{v\in\langle Y_{1},\dots,Y_{n-1},X_{1},\dots,X_{m}\rangle_{{\mathbb{F}_{p}}}}(Y_{n}+v),$
and then we can apply the identity (A.9) replacing $n$ with n1n-1 (satisfied by the inductive hypothesis) to show that (A.8) is equivalent to
$\Delta_{n}(P_{V_{m}}(Y_{1}),\cdots,P_{V_{m}}(Y_{n}))=\Delta_{n-1}(P_{V_{m}}(Y_{1}),\cdots,P_{V_{m}}(Y_{n-1}))\prod_{v\in\langle Y_{1},\dots,Y_{n-1},X_{1},\dots,X_{m}\rangle}(Y_{n}+v)$
@@
$\prod_{w\in\langle P_{V_{m}}(Y_{1}),\cdots,P_{V_{m}}(Y_{n-1})\rangle_{{\mathbb{F}_{p}}}}(P_{V_{m}}(Y_{n})+w)=\prod_{v\in\langle Y_{1},\dots,Y_{n-1},X_{1},\dots,X_{m}\rangle_{{\mathbb{F}_{p}}}}(Y_{n}+v).$
@@
When $m=1$ the identity of Lemma A.7 is already known. It appears for example in work of Elkies [Elkies99, p.81] and it is used in the proof of Proposition 5.3.
@@
Proposition A.11 (Proposition 2.3. of [FresnelMatignon23]).
Let $W\subset k$ be a ${\mathbb{F}}_{p}$ vector space of dimension $n$. Let $\underline{a}$ be an ${\mathbb{F}}_{p}$-basis of $W$ and $\underline{a}^{\star}=(a_{1}^{\star},\dots,a_{n}^{\star})$ be its dual basis. For every non-zero $n$-tuple $\underline{\epsilon}=(\epsilon_{1},\dots,\epsilon_{n})\in{\mathbb{F}}_{p}^{n}-\underline{0}$, we denote by $\varphi_{\underline{\epsilon}}$ the linear homomorphism $W\to{\mathbb{F}}_{p}$ given by $\varphi_{\underline{\epsilon}}:=\sum_{i=1}^{n}\epsilon_{i}a_{i}^{\star}$. Recall that we have defined $P_{\ker\varphi_{\underline{\epsilon}}}(X):=\prod_{w\in\ker\varphi_{\underline{\epsilon}}}(X-w)$.
@@
[…]
Proposition A.16 (Proposition 5.1. of [FresnelMatignon23]).
@@
Finally, we prove two results that are not in [FresnelMatignon23] and that are used in Section 4.
@@
Let $X_{1},\dots,X_{n}$ be variables over $\mathbb{F}_{p}$}, and for every $0\leq t\leq n$ let $V_{n-t}=\langle X_{1},\dots,X_{n-t}\rangle_{{\mathbb{F}}_{p}}$. Then, we have that
$\Delta_{n}(X_{1},\dots,X_{n})=\Delta_{n-t}(X_{1},\dots,X_{n-t})\Delta_{t}(P_{V_{n-t}}(X_{n-t+1}),\dots,P_{V_{n-t}}(X_{n})),$
@@
$\frac{\Delta_{n-1}(X_{1},\dots,\widehat{X_{i}},\dots,X_{n})}{\Delta_{n}(X_{1},\dots,X_{n})}=\frac{\Delta_{t-1}(P_{V_{n-t}}(X_{n-t+1}),\dots,\widehat{P_{V_{n-t}}(X_{i})},\dots,P_{V_{n-t}}(X_{n}))}{\Delta_{t}(P_{V_{n-t}}(X_{n-t+1}),\dots,P_{V_{n-t}}(X_{n}))}.$
@@
[…]